OpenStudy (anonymous):

Find the derivative with and without using the chain rule. f(x) = (x^3 - 1)^2

OpenStudy (anonymous):

Do you know how to expand it? \[ (x^3-1)^2 = (x^3-1)(x^3-1) \]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is that with the chain rule or?

OpenStudy (anonymous):

This is what is called "multiplication"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you factor it right?

OpenStudy (anonymous):

let me try to expand it

OpenStudy (anonymous):

Now we will use distributive property of multiplication. \[ (x^3-1)(x^3-1) = x^3(x^3-1)-1(x^3-1)=x^6-x^3-x^2+1=x^6-2x^3+1 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now take the derivative.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

the derivative of x^6 - 2x^3 + 1 right?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

i got 6x^2 (x^3 - 1)

OpenStudy (anonymous):

What is the derivative of \(x^6\)?

OpenStudy (anonymous):

What is the derivative of \(-2x^3\)?

OpenStudy (anonymous):

its 6x^5

OpenStudy (anonymous):

and -6x^2 for -2x^3

OpenStudy (anonymous):

so it will be 6x^5 - 6x^2 ?

OpenStudy (anonymous):

Yeah. You don't need to factor it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that's it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay can did we use the chain rule on this one or? no chain rule?

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

@wio ?

OpenStudy (sirm3d):

chain rule was not used

OpenStudy (anonymous):

okay thanks